# RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning

### v1

@@ 4   Methodology
In this section, we present a systematic framework for developing our proposed approach. We begin by integrating Distributional RL with Imitation Learning (IL). Next, we introduce a novel regularization technique that enhances existing Implicit Reward methods through an adaptive reward mechanism. Finally, we propose our algorithm, which leverages the strengths of both Distributional RL and IL to achieve superior performance in complex environments.
4.1   Distributional Value Function
Our method diverges from traditional imitation learning by modeling value distributions as critics in an actor-critic framework [Zhou et al., 2023]. We propose that learning the full return distribution $Z^{\pi}(s,a)$—rather than point estimates like $Q^{\pi}(s,a)$—enables better decision-making in complex environments. By optimizing policies using the expectation of $Z^{\pi}(s,a)$, we derive a more robust learning signal. This robustness stems from explicitly capturing environmental stochasticity through distributional modeling, avoiding the limitations of Q-networks that collapse variability into deterministic estimates. In Distributional RL, the action-value function is defined as $Q^{\pi}(s,a)=\mathbb{E}[Z^{\pi}(s,a)]$, preserving the mathematical properties required by the IQ-Learn framework, ensuring theoretical consistency (see Lemma A.1 for more details).
@@ 4.2   Implicit Reward Regularization
In this section, we propose a novel regularizer for Inverse Reinforcement Learning (IRL) that refines existing implicit reward formulations [Garg et al., 2021]. We define the implicit reward as $R_{Q}(s,a)=Q^{\pi}(s,a)-\gamma V^{\pi}(s^{\prime})$, where $V^{\pi}(s^{\prime})=Q^{\pi}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})$. Prior work often regularizes implicit rewards using $L2$-norms, treating them as squared Temporal Difference (TD) errors between rewards and fixed targets [Reddy et al., 2020, Al-Hafez et al., 2023]. While adopting a similar squared TD framework [Mnih et al., 2015], we introduce adaptive targets $\lambda^{\pi_{E}}$ (for expert policy $\pi_{E}$) and $\lambda^{\pi}$ (for imitation policy $\pi$) to derive our convex regularizer $\Gamma:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$:
$\Gamma(R_{Q},\lambda)=\mathbb{E}_{\rho_{E}}\left[(R_{Q}-\lambda^{\pi_{E}})^{2}\right]+\mathbb{E}_{\rho_{\pi}}\left[(R_{Q}-\lambda^{\pi})^{2}\right].$ (5)
Lemma 4.1.
The implicit reward structure $R_{Q}$ and regularizer $\Gamma$ (Equation 5) guarantee:
  • Bounded Rewards: $R_{Q}$ converges to a convex combination of adaptive targets $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$.
  • Temporal Consistency: Ensures $|Q(s,a)-\gamma Q(s^{\prime},a^{\prime})|\leq|\lambda|+|\epsilon|+\gamma\alpha|\log\pi(a^{\prime}|s^{\prime})|$,
    where $\epsilon$ is the optimization tolerance.
(Proof: Appendix A.)
This framework stabilizes learning by tethering rewards $R_{Q}$ to adaptive targets that evolve with the policy, unlike static methods (e.g., LSIQ/SQIL) that use fixed rewards (±$\pm 1$). By co-optimizing $R_{Q}$ and $\lambda$, we establish dynamic equilibrium—targets adapt to the current policy to avoid brittle regularization while maintaining bounded Q-values. The temporal consistency property ensures smooth value function updates, which directly translates to robust policy updates—reducing variance and improving the likelihood of convergence to the optimal policy. This dual adaptability distinguishes our approach from prior rigid regularization schemes.
4.3   Algorithm
We introduce RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning, an algorithm enhancing implicit reward IRL through a convex regularizer inspired by TD error minimization. Unlike standard RL that uses environmental rewards, RIZE employs self-updating adaptable targets, creating a somewhat reinforcement paradigm where rewards automatically align with their moving targets through regularization (see Algorithm 1).
Critic Updates
Let $Z_{\phi,\tau}^{\pi}(s,a)$ denote the return distribution and $\pi_{\theta}(a|s)$ the policy. We compute Q-values as:
$Q^{\pi}(s,a)=\sum_{i=0}^{N-1}(\tau_{i+1}-\tau_{i})Z_{\phi,\tau_{i}}^{\pi}(s,a)$ (6)
where $\tau_{i}$ are quantile fractions and $Z_{\phi,\tau_{i}}^{\pi}(s,a)$ represents quantile-specific return distributions.
The core objective integrates a squared-error regularizer $\Gamma(R_{Q},\lambda)$ (Equation 5) to stabilize target alignment:
$\mathcal{L}(\pi,Q)=$ $\mathbb{E}_{\rho_{E}}[R_{Q}]-\mathbb{E}_{\rho_{\pi}}[R_{Q}]-\mathcal{H}(\pi)$ (7)
$-c\left[\mathbb{E}_{\rho_{E}}[(R_{Q}-\lambda^{\pi_{E}})^{2}]+\mathbb{E}_{\rho_{\pi}}[(R_{Q}-\lambda^{\pi})^{2}]\right]$
where implicit rewards $R_{Q}$ derive from: $R_{Q}(s,a)=Q^{\pi}(s,a)-\gamma\left[Q^{\pi}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\right]$
Adaptive Target Updates
@@ 4.3   Algorithm
Targets $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ self-update via the following objectives, creating a feedback loop where reward estimates continuously adapt to match their moving targets.
$\underset{\lambda^{\pi_{E}}}{\min}\>\mathbb{E}_{\rho_{E}}\left[R_{Q}-\lambda^{\pi_{E}}\right]^{2},\quad\underset{\lambda^{\pi}}{\min}\>\mathbb{E}_{\rho_{\pi}}\left[R_{Q}-\lambda^{\pi}\right]^{2}$ (8)
Algorithm 1 Regularized Imitation Learning via Distributional Reinforcement Learning
1:  Initialize value distribution $Z_{\phi}$, policy $\pi_{\theta}$, and target rewards $\lambda^{\pi_{E}}$, $\lambda^{\pi}$
2:  for step $i$ in $\{1,\dots,N\}$ do
3:     Calculate $Q(s,a)=\mathbb{E}[Z_{\phi}(s,a)]$ using Equation 6
4:     Update value $Z_{\phi}$ using objective from Equation 7
5:
$\phi_{t+1}\leftarrow\phi_{t}-\beta_{Q}\nabla_{\phi}[-J(\phi)]$
6:     Update policy $\pi_{\theta}$ (like SAC)
7:
@@
$\theta_{t+1}\leftarrow\theta_{t}+\beta_{\pi}\nabla_{\theta}\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\theta}(\cdot|s)}[\min_{k=1,2}Q_{k}(s,a)-\alpha\log\pi_{\theta}(a|s)]$
8:     Update target rewards $\lambda^{\pi}$ and $\lambda^{\pi_{E}}$ using objectives in  8
9:
$\lambda^{\pi}_{t+1}\leftarrow\lambda^{\pi}_{t}-\beta_{\lambda}\nabla_{\lambda^{\pi}}\Gamma(R_{Q},\lambda)$
10:
$\lambda^{\pi_{E}}_{t+1}\leftarrow\lambda^{\pi_{E}}_{t}-\beta_{\lambda}\nabla_{\lambda^{\pi_{E}}}\Gamma(R_{Q},\lambda)$
@@ 5   Experiments
We evaluate our algorithm on four MuJoCo [Todorov et al., 2012] benchmarks (HalfCheetah-v2, Walker2d-v2, Ant-v2, Humanoid-v2) against state-of-the-art imitation learning methods: IQ-Learn [Garg et al., 2021], LSIQ [Al-Hafez et al., 2023], and SQIL [Reddy et al., 2020]. All experiments are conducted with five seeds for statistical significance [Henderson et al., 2018]. To assess sample efficiency, we test each method using three and ten expert trajectories. For ten trajectories, we retain baseline hyperparameters; for three trajectories (not reported in prior works), we adapt configurations from single-demonstration settings. IQ-Learn and LSIQ use their official implementations, while SQIL is evaluated using the LSIQ codebase. Results report mean performance across five seeds, with half a standard deviation to indicate uncertainty, and lines are smoothed for better readability. We normalize episode returns based on expert performance.
Figure 1: Normalized return of RIZE vs. online imitation learning baselines on Gym MuJoCo tasks. We depict the sorted top 25% episodic returns across five seeds to evaluate convergence to expert-level behavior. We evaluate with three and ten expert trajectories.
@@ 5.1   Main Results
Our method outperforms LSIQ and SQIL across tasks, with IQ-Learn as the only competitive baseline. Notably, in the most complex environment, Humanoid-v2, our approach is the sole method achieving expert-level performance, while all baselines fail (see Figure 1). This demonstrates our algorithm’s scalability to high-dimensional control problems. Additionally, our method shows superior sample efficiency, requiring fewer gradient steps to match expert performance compared to SOTA algorithms. These results highlight the robustness and efficiency of our approach in tackling complex tasks (see Figure 2).
Figure 2: Normalized return of RIZE vs. online imitation learning baselines on Gym MuJoCo tasks. We use 10 expert trajectories for all tasks.
5.2   Reward Dynamics Analysis
Our implicit reward regularizer $\Gamma(R_{Q},\lambda)$ (Equation 5), ensures proximity between the learned $R_{Q}$ and target rewards $\lambda$. As illustrated in Figure 3, when provided with sufficient expert trajectories (ten demonstrations), the rewards for both the expert and the policy stabilize around consistent values, as observed in environments like HalfCheetah and Ant. While our method is not Adversarial IL, it extends MaxEnt IRL principles. This alignment is evident as the policy increasingly mirrors expert behavior, causing the discriminator to converge and provide similar reward signals for both the expert and policy samples. This convergence reflects an equilibrium between policy optimization and reward learning, demonstrating the stability and effectiveness of our approach.
Figure 3: Implicit reward curves for expert and policy samples on Gym MuJoCo tasks. We use 10 expert trajectories for all tasks.
@@ 5.3   Ablation: Regularization Strategies
We investigate the effect of squared TD error regularization on the Ant-v2 task using three distinct formulations: Expert-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}-\lambda^{\pi_{E}}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$, Policy-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}-\lambda^{\pi}]^{2}$, Baseline L2: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$.
Our results in Figure 4 demonstrate that applying squared TD regularization to both expert and policy samples improves robustness. While the expert-focused TD formulation outperforms the policy-focused variant, neither consistently achieves expert-level behavior. Over time, both exhibit gradual performance degradation, underscoring the challenges of maintaining stable imitation. This highlights the need for more refined regularization strategies to ensure consistent and reliable learning in imitation learning.
Figure 4: Ablation study on the effect of various regularization strategies on the Ant-v2 task, evaluated using three demonstrations. Turquoise represents RIZE using the convex regularizer $\Gamma$.

### v2

@@ 4 Methodology
This section introduces a framework that integrates Distributional Reinforcement Learning with Inverse RL. We first show how value distributions can replace point-estimate critics. We then propose an adaptive regularization technique for implicit rewards and analyze its properties. Finally, we derive RIZE, which combines distributional critics with bounded-reward imitation learning.
4.1 Distributional Value Integration
Our approach departs from traditional imitation learning by explicitly using value distributions as critics within an actor-critic framework (Zhou et al., 2023). We argue that learning the soft value distribution $Z(s,a)$ in Equation (1), rather than relying solely on point estimates like $Q(s,a)$, enhances decision-making by capturing uncertainty in complex environments. This view is consistent with recent neuroscience findings suggesting that decision-making in the prefrontal cortex relies on learning distributions over outcomes rather than only their expectations (Muller et al., 2024).
Moreover, access to the full return distribution enables the computation of statistical moments—most notably the expectation—which we optimize both policy and critic using the expectation of $Z$ (Bellemare et al., 2017; Dabney et al., 2018b; a), yielding a more robust learning signal while remaining compatible with IQ-Learn.
We compute $Q$ as the expectation of the soft value distribution:
$Q(s,a)=\sum_{i=0}^{N-1}(\tau_{i+1}-\tau_{i})\,Z_{\tau_{i}}(s,a)\,$, (5)
where $\{\tau_{i}\}$ are quantile fractions and $Z_{\tau_{i}}(s,a)$ denotes the corresponding quantile values (see Lemma A.1).
@@ 4.2 Implicit Reward Regularization
In this section, we propose a regularizer for inverse RL that refines existing implicit reward formulations (Garg et al., 2021). The implicit reward is defined as:
$R_{Q}(s,a)=Q(s,a)-\gamma\mathbb{E}_{P,\pi}\left[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\right].$ (6)
@@
Previous works typically regularize implicit rewards either using $L_{2}$-norms (Garg et al., 2021) or by treating them as squared-TD errors between rewards and fixed targets (Reddy et al., 2020; Al-Hafez et al., 2023). While we adopt a similar squared-TD setting, we introduce adaptive targets $\lambda^{\pi_{E}}$ (for the expert $\pi_{E}$) and $\lambda^{\pi}$ (for the imitation policy $\pi$) to construct our convex regularizer $\Gamma:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$:
$\Gamma(R_{Q},\lambda)=\mathbb{E}_{\rho_{E}}\left[(R_{Q}(s,a)-\lambda^{\pi_{E}})^{2}\right]+\mathbb{E}_{\rho_{\pi}}\left[(R_{Q}(s,a)-\lambda^{\pi})^{2}\right].$ (7)
@@
These targets self-update through a feedback loop where reward estimates continuously adapt to match moving targets:
$\underset{\lambda^{\pi_{E}}}{\min}\>\mathbb{E}_{\rho_{E}}\left[\left(R_{Q}(s,a)-\lambda^{\pi_{E}}\right)^{2}\right]\,,\qquad\underset{\lambda^{\pi}}{\min}\>\mathbb{E}_{\rho_{\pi}}\left[\left(R_{Q}(s,a)-\lambda^{\pi}\right)^{2}\right]\,.$ (8)
Substituting $\Gamma(R_{Q},\lambda)$ for the $L_{2}$ term in Equation (4) and using Equation (5) to compute $Q$, we obtain:
$\mathcal{L}(\pi,Q)$ $=\mathbb{E}_{\rho_{E}}[R_{Q}(s,a)]-\mathbb{E}_{\rho_{\pi}}[R_{Q}(s,a)]-\alpha\mathcal{H}(\pi)$ (9)
$\qquad-c\Big{[}\mathbb{E}_{\rho_{E}}\left[(R_{Q}(s,a)-\lambda^{\pi_{E}})^{2}\right]+\mathbb{E}_{\rho_{\pi}}\left[(R_{Q}(s,a)-\lambda^{\pi})^{2}\right]\Big{]}\,$,
where $c$ is the regularization coefficient.
As in IQ-Learn and LSIQ, our method seeks behavior indistinguishable from expert demonstrations in a non-adversarial implicit-reward setting. Prior work analyzes optimal implicit rewards, drawing on Max–Min analyses from GANs (Al-Hafez et al., 2023; Goodfellow et al., 2014). Because we bind rewards to adaptive targets via $\Gamma(R_{Q},\lambda)$ (Equation (7)), we carry out an analogous analysis, stated below.
Proposition 4.1.
Let $R_{Q}(s,a)=(\mathcal{T}^{\pi}Q)(s,a)$ denote the implicit reward derived from point-estimate Q-values, where $Q(s,a)=\mathbb{E}[Z(s,a)]$. Let $\rho_{E}(s,a)$ and $\rho_{\pi}(s,a)$ denote occupancy measures under $\pi_{E}$ and $\pi$, respectively. For fixed $\pi$, the optimal TD-regularized reward satisfies:
$R_{Q}^{*}(s,a)=\,\frac{\rho_{E}(s,a)-\rho_{\pi}(s,a)}{(2c)\,(\rho_{E}(s,a)+\rho_{\pi}(s,a))}+\frac{\rho_{E}(s,a)\lambda^{\pi_{E}}+\rho_{\pi}(s,a)\lambda^{\pi}}{\rho_{E}(s,a)+\rho_{\pi}(s,a)}\,.$ (10)
Proof.
Differentiating $\mathcal{L}(\pi,Q)$ in Equation (9) with respect to $R_{Q}(s,a)$ (holding $\pi$ fixed) and setting the result to zero yields
$0=\rho_{E}-\rho_{\pi}-2c\big{[}\rho_{E}(R_{Q}-\lambda^{\pi_{E}})+\rho_{\pi}(R_{Q}-\lambda^{\pi})\big{]}\,$, (11)
from which Equation (10) follows. ∎
As established in Corollary A.2, this optimal reward satisfies:
$R_{Q}^{*}(s,a)\in\left[-\frac{1}{2c}+\lambda_{\min},\;\frac{1}{2c}+\lambda_{\max}\right]\,$, (12)
where $\lambda_{\min}:=\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$ and $\lambda_{\max}:=\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$. The coefficient $c$ and targets bound rewards within a well-defined range. In practice, we use $\lambda\in[0,10]$ and $c\in[0.1,0.5]$, which we found to promote stable training (see Appendix B.3).
Moreover, when the occupancy measures match, the optimal reward simplifies significantly, as stated in Corollary A.3:
$R_{Q}^{*}(s,a)=\lambda^{\pi_{E}}=\lambda^{\pi}\,.$ (13)
Recalling that we represent rewards through Q-values, the boundedness of the optimal reward ensures that critic updates remain constrained and—because the policy directly depends on these critic values—promotes stable policy optimization. However, as previously noted (Al-Hafez et al., 2023; Viano et al., 2022), the convergence guarantee originally stated for IQ-Learn (Garg et al., 2021) does not extend to the $\chi^{2}$-regularizer used in practice. Consequently, a formal proof for the resulting alternating SAC updates remains an open question and is left for future work.
4.3 Practical Algorithm
We now introduce RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning (Algorithm 1). We approximate $Z$ and $\pi$ with neural networks, optimizing via $Q=\mathbb{E}[Z]$. Following Distributional SAC (Ma et al., 2020), we employ target policies for next-action sampling and a double-critic architecture with target networks for stability. Lower learning rates proved critical for robust policy updates, with a four-layer MLP policy network necessary for complex tasks.
Targets $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ are optimized to match expected rewards. We initialize $\lambda^{\pi_{E}}$ higher than $\lambda^{\pi}$ and update $\lambda^{\pi}$ with lower learning rates due to its sensitivity.
Algorithm 1 RIZE
1:Initialize $Z_{\phi}$, $\pi_{\theta}$, $\lambda^{\pi_{E}}$, and $\lambda^{\pi}$
2:for step $t$ in $\{1,\dots,N\}$ do
3:  Calculate $Q(s,a)=\mathbb{E}[Z_{\phi}(s,a)]$ using Eq. 5
4:  Update $Z_{\phi}$ using Eq. 9
5:  $\phi_{t+1}\leftarrow\phi_{t}-\beta_{Z}\nabla_{\phi}[-\mathcal{L}(\phi)]$
6:  Update $\pi_{\theta}$ (like SAC)
@@ 4.3 Practical Algorithm
7:  $\begin{aligned} \theta_{t+1}\leftarrow\theta_{t}+\beta_{\pi}\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}s\sim\mathcal{D},\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}[&\min_{k=1,2}Q_{k}(s,a)\\ &-\alpha\log\pi_{\theta}(a|s)]\end{aligned}$
8:  Update $\lambda^{\pi}$ and $\lambda^{\pi_{E}}$ using Eq. 8
9:  $\lambda^{\pi}_{t+1}\leftarrow\lambda^{\pi}_{t}-\beta_{\lambda^{\pi}}\nabla_{\lambda^{\pi}}\Gamma(R_{Q},\lambda)$
10:  $\lambda^{\pi_{E}}_{t+1}\leftarrow\lambda^{\pi_{E}}_{t}-\beta_{\lambda^{\pi_{E}}}\nabla_{\lambda^{\pi_{E}}}\Gamma(R_{Q},\lambda)$
@@ 5 Experiments
We evaluate our algorithm on four MuJoCo (Todorov et al., 2012) benchmarks (HalfCheetah-v2, Walker2d-v2, Ant-v2, Humanoid-v2) against state-of-the-art imitation learning methods: IQ-Learn (Garg et al., 2021), LSIQ (Al-Hafez et al., 2023), and SQIL (Reddy et al., 2020). All experiments are conducted with five seeds for statistical significance (Henderson et al., 2018). We assess each method using three and ten expert trajectories. For ten trajectories, we retain baseline hyperparameters; for three trajectories (not reported in prior works), we adapt configurations from single-demonstration settings. IQ-Learn and LSIQ use their official implementations, while SQIL is evaluated using the LSIQ codebase. Results report mean performance across five seeds, with half a standard deviation to indicate uncertainty, and lines are smoothed for better readability. We normalize episode returns based on expert performance. For details, expanded visualizations, and hyperparameter tuning, see Appendix B.
Figure 1: Normalized return of RIZE versus online imitation learning baselines on Gym MuJoCo tasks. We depict the sorted top 25% episodic returns across five seeds to evaluate convergence to expert-level performance, using three and ten expert trajectories.
@@ 5.1 Main Results
Our method outperforms LSIQ and SQIL across tasks, with IQ-Learn as the only competitive baseline. Notably, in the most complex environment, Humanoid-v2, our approach is the sole method achieving expert-level performance, while all baselines fail (see Figures 1, 2). This demonstrates our algorithm’s scalability to high-dimensional control problems. Additionally, our method shows superior sample efficiency, requiring fewer gradient steps to match expert performance compared to SOTA algorithms. These results highlight the robustness and efficiency of our approach in tackling complex tasks.
5.2 Recovered Reward
Here, we compare the reward behavior of RIZE against the baselines. As shown in Figure 3, our approach effectively bounds the rewards for both expert and policy samples across environments. The reward regularizer $\Gamma(R_{Q},\lambda)$ keeps the recovered rewards close to adaptive targets, with the optimization of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ via Equation (8) enforcing an upper bound on expert rewards and a lower bound on policy rewards. With sufficient expert trajectories, rewards for both expert and policy samples stabilize around consistent values. In contrast, baseline methods exhibit unbounded reward growth in the Humanoid task, and SQIL fails to constrain rewards even in the simpler HalfCheetah environment.
Figure 3: Implicit reward curves for expert and policy samples on MuJoCo tasks. We evaluate on ten expert demonstrations.
5.3 Role of Distributional Value Function
We examine the evolution of critics in RIZE compared to the baselines. Unlike the baselines, which train Q-networks, RIZE employs distributional value functions and uses their expectation $\mathbb{E}[Z(s,a)]$ for loss optimization. Since all methods update policies by maximizing state-action values, unbounded values can significantly undermine robustness. As illustrated in Figure 4, RIZE and LSIQ maintain bounded state-action values in the challenging Humanoid environment, whereas SQIL and IQ-Learn do not. Nevertheless, the boundedness of critics alone does not guarantee expert-like policy performance, as evidenced by LSIQ’s results in Figure 2.
Figure 4: State-action value curves for expert and policy samples on MuJoCo tasks. Ten trajectories are used throughout the analysis.
Moreover, to assess the significance of quantile-based value networks Z(s,a)\,Z(s,a), we train RIZE with point-estimate critics Q(s,a)\,Q(s,a) and compare the results with the original implementation. As shown in Figure 5, distributional value functions consistently outperform Q-networks across all tasks, exhibiting lower variance and greater sample efficiency. Notably, in the Humanoid environment, Q-values fail to replicate expert demonstrations, whereas quantile-based critics maintain expert-level performance.
Figure 5: Normalized return of RIZE on Gym MuJoCo tasks utilizing value distribution $Z(s,a)$ or classic $Q(s,a)$ as the critic. We use three expert trajectories for all tasks.
@@ 5.4 Ablation: Regularization Strategies
We investigate the effect of squared-TD error regularization on the Ant-v2 task using three distinct formulations: Expert-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}-\lambda^{\pi_{E}}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$, Policy-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}-\lambda^{\pi}]^{2}$, Baseline L2: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$.
Our results in Figure 6 demonstrate that applying squared-TD regularization to both expert and policy samples improves robustness. While the expert-focused TD formulation outperforms the policy-focused variant, neither consistently achieves expert-level behavior. Over time, both exhibit gradual performance degradation, underscoring the challenges of maintaining stable imitation. This highlights the need for more refined regularization strategies to ensure consistent and reliable learning in imitation learning.
Figure 6: Ablation study on the effect of various regularization strategies, evaluated using three demonstrations.